OpenStudy (anonymous):

NEEDED ASAP PLEASE HELP !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!Solve for x: 2 over 5 (x - 4) = 2x. -8 negative 1 over 2 1 -1

OpenStudy (anonymous):

the listed is the possible answer

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 (x-4)} ~ \color{red}{or}~ \frac{ 2 }{ 5 } (x-4)\]

OpenStudy (mertsj):

\[\frac{2}{5(x-4)}=\frac{2x}{1}\]

OpenStudy (mertsj):

Cross multiply and solve

OpenStudy (anonymous):

sorry what i wrote is wrong its 2 over 5(x-4)=2x

OpenStudy (mertsj):

Does it look like this: \[\frac{2}{5}(x-4)=2x\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

1. distribute the 2/5 2. Subtract 2/5 x from both sides 3. Multiply both sides by 5/8

OpenStudy (anonymous):

how do i distribute that this is where i got confused

OpenStudy (mertsj):

3(2x-8)=6x-24 In other words, multiply everything inside the parentheses by what is outside.

OpenStudy (anonymous):

waaaaaaaaaaat?????????????????????/

OpenStudy (mertsj):

Do you know how to multiply?

OpenStudy (anonymous):

where did that equation come from? @Mertsj

OpenStudy (mertsj):

I made it up to show you how to distribute

OpenStudy (mertsj):

\[\frac{2}{5}(x-4)\] You need to distribute the 2/5

OpenStudy (anonymous):

yes but its a fraction i cant put 2/5 into 4 evenly

OpenStudy (mertsj):

Sometimes when you multiply by a fraction, you end up with a fraction.

OpenStudy (anonymous):

oh so just by that is it safe to say the answer is -1/2

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

no

OpenStudy (anonymous):

oh

OpenStudy (mertsj):

Why don't you just go ahead and solve it? What's the big deal?

OpenStudy (anonymous):

no big deal i am solving it i just thought if it turns out a fraction maybe the answer is a fraction